Circular No. 8998

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

1998 WV_{24}

S. D. Benecchi, K. S. Noll, and E. A. Barker, Space Telescope Science Institute; W. M. Grundy, Lowell Observatory; and H. F. Levison, Southwest Research Institute, report that the transneptunian object 1998 WV_{24} (cf. MPECs 1998-X12 and 1999-V06) is a binary. The observations were made during 2007 Aug. 27.1069–27.1375 UT with the Planetary Camera of the Wide Field Planetary Camera 2 on the Hubble Space Telescope, using the F606W filter (wide V) with one 260-s exposure at four dithered positions on the detector. The two components were separated by an angular distance of 0′′.051 ± 0′′.002, with the secondary fainter by 0.3 magnitude. The secondary was located at 0′′.033 ± 0′′.003 in α and 0′′.039 ± 0′′.003 in δ relative to the primary. Component centroids were determined by fitting the point-spread function.

V2468 CYGNI D. K. Lynch, R. W. Russell, and R. J. Rudy, The Aerospace Corpora-

tion; M. Sitko, Space Science Institute; and C. E. Woodward, University of Minnesota, report 0.8- to 5.5-μm spectroscopy of the nova V2468 Cyg (cf. IAUC 8927, 8928, 8936) on Oct. 4 UT using the Infrared Telescope Facility (+ Spex). Though V2468 Cyg had faded by more than a factor of two since mid-August, the coronal lines have strengthened significantly. The [Si VI] line at 1.96 μm is among the strongest of the emission features. Lines of [S VIII], [S IX], [Al IX], and [Ca VIII] are also present. Several Rydberg lines are present, including those at 0.89, 1.19, 1.11, 1.55, and 2.1 μm.

V2670 OPHIUCHI M. Sitko, Space Science Institute; D. K. Lynch, R. W. Russell, and R.

J. Rudy, The Aerospace Corporation; and C. E. Woodward, University of Minnesota, report 0.8- to 5.5-μm spectroscopy of the nova V2670 Oph (cf. IAUC 88947, 8948, 8949, 8956) on Oct. 4 UT using the Infrared Telescope Facility (+ Spex). The spectrum remains low in excitation, still displaying emission lines of Fe II, N I, and O I. However, the C I lines that were strong in mid-June are barely detectable, and the He I 1.083-μm line is now the strongest feature in the infrared spectrum. The broad components of the emission lines have disappeared entirely. This nova has not formed dust and is unlikely to do so at this late stage.

2008 October 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green